# Phase field crystal modelling of deformation assisted precipitation in binary alloys

### v1

@@ Abstract
We consider the process of precipitation in binary alloys in the presence of mechanical deformation. It is commonly observed that mechanical deformation prior to or during precipitation leads to microstructure with excess defects, which allows for enhanced precipitate nucleation and growth rates [1, 2, 3]. To investigate this phenomenon, we employ a two-dimensional phase-field crystal alloy model endowed with a temperature dependent mobility, making it capable of recovering isothermal transformation (TTT) diagrams with a characteristic inflection point (nose) about a critical temperature. We examine the variation in the time-scale of precipitation and its connection to the time-scale of the applied deformation, focusing on the roles of atomic defects in the processes involved. Our results indicate that precipitation is initially delayed through application of a deformation until a critical strain is achieved, beyond which precipitation proceeds more rapidly, assisted by plastic deformation such as grain boundary buckling or dislocation nucleation. We show that the evolution of the precipitated fraction, $f(t)$, departs from classical Avrami behaviour. Specifically, $df/dt$ develops two peaks indicative of a “plateau"-like inflection in $f(t)$, signalling the transition to defect assisted precipitate nucleation. We analyze these plateaus as a function of the deformation rate and demonstrate that the they exhibit a discontinuous bifurcation as the time-scale of applied deformation is increased. These findings are compared to and found to be consistent with experiments.
@@
Mechanical deformations driving out-of-equilibrium behaviours are ubiquitous in current and emerging material processing techniques such as additive manufacturing, laser shock peening, friction stir-welding, etc. [4, 5, 6]. The material conditions left behind in the wake of these processes provide inflated driving forces for diffusive and phononic relaxation due to increased saturation of solid solutions and the presence of point defects. This is in sharp contrast to classical casting conditions [7, 8]. An emphasis on the physics governing the aforementioned material processes has thus been of interest to the materials science community, particularly how such processes deviate from classically understood mechanisms of precipitation, recrystallization and recovery in metal alloys.
Precipitation readily occurs at sites of higher energy, such as defects and grain boundaries. Mechanical deformation of crystalline solids produces many such sites, thus modifying the nucleation, growth, and coarsening properties encountered [3]. Precipitation can thus be seen as competitive with other forms of strain relaxation eg. recrystallization and recovery in metals and their alloys. These processes all use stored driving forces to relieve stress accruing at high energy sites within a deformed sample. Indeed, thus has been studied previously [1, 2], where experiments and numerical simulations have examined modifications to isothermal transformation (TTT) diagrams in the presence of mechanical deformation. Analytical models have likewise been formulated to quantify this interactive growth process [9]. These models typically require a priori knowledge of specific material properties which are difficult to extract such as activation energies, defect fractions. Furthermore, these models require knowledge of the nucleation and growth mechanisms of both recrystallized and precipitated grains.
@@
In order to investigate the atomistic behaviour of precipitation in the presence of deformation, this work examines a prototypical scenario of an equal tilt bicrystal undergoing a continuous compressive deformation with strain rate $\dot{\epsilon}$. To this end, we employ a structural phase field crystal (XPFC) model with a free energy functional that models substitutional binary alloys and has previously been applied to studies involving precipitation [22, 23], as well as deformation in such systems [24]. Our free energy is formulated as follows,
@@
where $c_{0}$ is the concentration at the reference. The final, excess, term of the free energy controls the length scale and strength of atomic-scale interactions. The correlation function $C_{nn}$ interpolates between pure $A$ and pure $B$ structures through interpolation functions $\chi_{A}(c)=1-3c^{2}+2c^{3}$, $\chi_{B}(c)=1-3(1-c)^{2}+2(1-c)^{3}$ such that $C_{nn}(\mathbf{r}-\mathbf{r^{\prime}})=\chi_{A}C_{AA}^{(2)}(\mathbf{r}-\mathbf{r^{\prime}})+\chi_{B}C_{BB}^{(2)}(\mathbf{r}-\mathbf{r^{\prime}})$. In XPFC, the pure species correlation functions are defined in Fourier space (with wave-vector magnitude $k$) by a sum of Gaussian functions centered about the dominant reciprocal lattice wave-vector ($k=k_{i}$) and $k=0$ modes of each species respectively,
@@
where $i=A,B$. The parameter $T_{i}^{0}$ controls the temperature scaling of the corresponding correlation function, $\sigma_{i,0}$ and $\sigma_{i,1}$ control the Gaussian widths of the $k=0$ and dominant wave-vector respectively. Finally, $A_{i,0}$ is constant that controls the amplitude of the $k=0$ mode. In this letter, we work in two spatial dimensions and employ a single wave-vector in our correlation to stabilize a hexagonal structure in each solid phase.
@@
where $\mu_{n/c}$ are the chemical potentials, $M_{n/c}$ are the mobilities and $\xi_{n/c}$ are noise sources applied to the $n$ and $c$ fields respectively. We also add an additional time-scale for quasi-phonon relaxation into the density equation. A full two-time scale model involving the couplings to $c$ are outside of the scope of this work. The noise terms satisfy their respective fluctuation dissipation relations, $\langle\xi_{n/c}(\textbf{r},t)\xi_{n/c}(\textbf{r'},t^{\prime})\rangle=N_{n/c}M_{n/c}k_{B}T\delta(\textbf{r}-\textbf{r'})\delta(t-t^{\prime})$ where $N_{n/c}$ subsume remaining constant prefactors and are treated as model parameters. We note that these dynamics are not strictly consistent in the regime where the average density of the system has significant variations [25]. A more rigorous albeit more computationally expensive treatment of the dynamics has been implemented in [26]. However, as the common tangents between the free energy wells of our system exist at small and nearly identical average densities, this approximation is justified.
While our systems are assumed to be isothermal, We expand on the model by introducing temperature dependent mobilities into the system dynamics. As will be shown below, this allows for the generation of TTT curves that can be controlled through model parameters. An Arrhenius form is assumed for the mobilities of both $n$ and $c$,
@@
where the activation energies $E_{A}^{n}$, $E_{A}^{c}$ are model parameters which control the curvature and position of the inflection point of a given TTT curve for precipitation. Strictly speaking, these will be coupled to the mobilities of the individual atomic species in a non-linear manner. For convenience, we treat these mobilities and activation energies as independent model parameters for the purposes of this study. It is also noted that the modified mobilities will shift the effective length-scale over which quasiphonons propagate undamped for fixed values of $\beta$.
@@
Figure 1: The isothermal transformation (TTT) diagram of precipitation at zero strain rate. The three sets of curves represent different fractions of completed transformation. A characteristic inflection or nose shape can be observed over the range of temperatures selected as a result of the temperature dependent mobilities defined in Eq. (6). Curves are plotted for a selection of $E_{A}^{c}$, which monotonically controls the reaction rate as shown.
@@
Figure 2: The sequence of events of a simulation of strain enhanced precipitation. Strain is applied along the $y$-direction. Panels ($\text{a}_{1}$)-($\text{a}_{4}$) show the evolution of the composition field and panel $\text{a}_{5}$ shows the associated evolution of the precipitated fraction $f$ as well as $df/dt$ in black and red, respectively, with the values corresponding to panels ($\text{a}_{1}$)-($\text{a}_{4}$) labelled with circles. Panels ($\text{b}_{1}$)-($\text{b}_{4}$) show the behaviour of the local orientation field and ($\text{b}_{5}$) shows the histogram of orientations of panels ($\text{b}_{1}$) and ($\text{a}_{4}$). The first peak in ${df}/{dt}$, corresponding to panels ($\text{a}_{2}$), $\text{b}_{2}$ is associated with the precipitation on the buckled interface. The second peak, corresponding to panels $\text{a}_{3}$, $\text{b}_{3}$ coincides with precipitate nucleation on dislocations. Panel $\text{b}_{4}$ shows the late stage coarsening behaviour of the system, which includes the reorientation of A-rich regions close to the grain boundary down to orientations intermediate between zero and the initial orientation of ±$\pm 6^{\circ}$.
@@
For all simulations described, we performed a quench below the solvus line into solid-solid coexistence. Full simulation parameters are presented in the Supplemental Material. We initialize the density field of the crystalline regions by seeding in an equal tilt bicrystal through a one-mode expansion. The lattice constant of this solid solution is set to be commensurate to the value attained by the interpolated correlation function for the given choice of supersaturation. The system is set to be initially strain-free by performing a minimization over the angle and number of atoms in the x and y directions for one (and by symmetry the other) grain as in [27]. The initial misorientation during this minimization was kept within the range 12.00±$\pm 0.05^{\circ}$. The average density and supersaturation are set to be initially uniform and equal to 0.05 and 0.15, respectively.
The density field is next allowed to relax by setting $M_{c}$ to zero. This is done so that the amplitude of the density can equilibrate prior to the application of a strain. Then both $n$ and $c$ are subsequently evolved simultaneously by Eqs. 4 and 5, with the application of a biaxial deformation with the tensile direction parallel to the grain boundary.
@@
In order to characterize the behaviour of the system over the course of simulations, three quantities are tracked: The grand potential energy, precipitated fraction $f$, and local orientation. The grand potential will be used to evaluate the dimensionless system pressure $\bar{P}=P/k_{B}T^{0}\rho^{0}V$ as in [28]. The orientation is extracted through local convolutions with oriented density wavelets as in [29]. Finally to calculate the precipitated fraction, local thresholding in composition is applied in conjunction with a Fourier space mask sharply peaked around the dominant wave-vector of the precipitating phase.
We first examine the situation where $\dot{\epsilon}=0$. While precipitation has been studied with XPFC in the past, to our knowledge no study has been performed with temperature variations in the mobility such as the ones proposed in Eq. (6). We employ the same initial conditions in $n$ and $c$ over a range of values of $T$ and $E_{A}^{c}$ and monitor the precipitate fraction, $f$ over the course of dynamical evolution of Eqs. (4) and (5). The fraction $f$ is calculated as described in the previous paragraph, with the compositional filtering chosen at a threshold compositional value of the average of the initial alloy composition and equilibrium B composition values.
@@
Over the course of each isothermal simulation, fluctuations drive the formation of precipitates along the grain boundary (GB) of the system. A detailed analysis of the geometry, growth rate and nucleation rate of these precipitates is outside of the purview of this paper. Over all temperatures and activation energies examined, the precipitate fraction obeys an Avrami-like behaviour
@@
where $m$ is an exponent which is determined by the dimensionality and mechanism of the transformation, $f^{eq}$ is the equilibrium precipitate fraction and $\tau_{0.5}$ is a characteristic time-of the system which is a function of the nucleation and growth-rate. The system pressure was found to exhibit a single dominant peak corresponding to the end of precipitate nucleation. Beyond this time, the system reduces its free energy and pressure through diffusion of solute out of the surrounding supersaturated bulk and into precipitate growth.
Figure 1 shows the isothermal transformation diagram (TTT) curve for a range of model temperatures, demonstrating the inflection about a critical temperature. The scaling of this surface is controlled by the usual driving force for precipitation (competition between bulk and surface energies) as well as the new mobility used here, which also affects the noise strength. At each temperature, the bulk energies are determined through the curvatures of the free energy landscape about the equilibrium wells of Eq. (1) in the space of average density, supersaturation and amplitude of the atomic density field $n$. The surface energy will be dependent on all of the above along with the effective gradient coefficient of the amplitude arising from the excess energy term, as well as $W_{c}$. These parameters are all held fixed for each value of $T$ along the curves shown in Figure 1.
@@
The evolution of a precipitating system under mechanical deformation is shown in Figure 2. At lower values of $\dot{\epsilon}$, $f$ obeys the same qualitative Avrami behavior as the undeformed case. Prior to precipitation, in order to compensate for the applied deformation, the density field undergoes elastic accommodation. The characteristic time-scale, $\tau_{0.5}$, for the precipitation is observed to initially increase with $\dot{\epsilon}$. This occurs due to an accelerated rotation of the bicrystal with increasing $\dot{\epsilon}$, causing a reduction in the effective driving force for precipitation along the boundary. The increase we observe in $\tau_{0.5}$ appears to qualitatively follow a simple $\exp(\gamma^{2})$ dependence in the time-scale for precipitation where $\gamma$ is the bicrystal surface energy’s misorientation dependent component. It is noted that the XPFC model has been shown to reproduce the correct Read-Shockley scaling of solute segregation and grain boundary energy as a function of misorientation in prior studies [30]. We have also performed simulations in the undeformed case with varying orientation that follow this same scaling.
@@
Figure 3 plots the system pressure as a function of time at different strain rates $\dot{\epsilon}$. As in the zero deformation case, at low $\dot{\epsilon}$ the pressure is first found to exhibit a single dominant peak at the initial stages of precipitation, followed by oscillations indicating accommodations by the grain boundary and the formation of any dislocations necessary to support the growth of the population of precipitates.
@@
Figure 3: The system pressure is plotted as a function of system time for different choices of the strain rate. At lower values of $\dot{\epsilon}$ (deep blue), an initial peak followed by oscillations indicating nucleation of dislocations post-precipitation is observed. At larger $\dot{\epsilon}$ (blue–cyan range) a second peak emerges corresponding to a new dislocation assisted precipitation mechanism. At the largest range of $\dot{\epsilon}$ simulated (yellow-red range), a single peak emerges again indicating precipitation assisted through grain boundary buckling.
Figure 4: The times when $f/f^{eq}$ attains $0.05,0.50$ and $0.95$ plotted against strain-rate. Also shown in blue is the time at which the system achieves its maximum in pressure $t_{P(max)}$ at each strain-rate. At low $\dot{\epsilon}$, $t_{P(max)}$ corresponds to the initialization of precipitation. At a critical $\dot{\epsilon}=2\times 10^{-6}$, $t_{P(max)}$ diverges significantly from its value during the initiation of precipitation. This occurs because the prominence of the second, plasticity induced, pressure peak surpasses the peak associated with precipitation, as shown in Figure 3. Upon further increase of $\dot{\epsilon}$ beyond $5\times 10^{-6}$, buckling of the GB occurs and $t_{P(max)}$ monotonically decreases past its initial value.
@@
As strain rate is increased further beyond a critical strain rate of $2\times 10^{-6}$, we observe systematic deviation in the precipitation mechanism. We found three mechanisms that contribute to this deviation, each activating in order of increasing $\dot{\epsilon}$.
@@
At the lower range of $\dot{\epsilon}$, we observe the nucleation of dislocations from initial precipitates or the GB, also shown in Figure 2. It is well known that dislocations act as preferential nucleation sites for precipitation [31]. This has also been examined numerically in prior studies using the XPFC model [22, 23]. If these dislocations nucleate sufficiently quickly, excess solute may diffuse to not only the GB but also towards some dislocations flowing into the bulk of the system. The diffusion of solute towards these dislocations can then induce further precipitation at sites away from the grain boundary. These dislocation nucleation events are characterized by the formation of a new peak in the system pressure that progressively increases in prominence and occurs after initial precipitation along the GB. This peak is subsumed into the pressure shift due to precipitation at intermediate strain rates. This effect is also seen in the data of Figure 3 and Figure 4. The same effect also leads to the emergence of a second peak in $df/dt$ as shown in Figure 5, corresponding to additional precipitate nucleation mechanism.
@@
Figure 5: Panel (a) shows the precipitated fractions (green) and their respective derivatives (black), indicating the first and second peaks (blue and red respectively). The data are plotted for three values of $\dot{\epsilon}$, at progressively higher deformation strain rates. As can be seen the low strain-rate scenario demonstrates a single peak whereas the second and third cases shown here exhibit a prominent second peak in $df/dt$, associated with dislocation nucleation. Panel (b) shows the corresponding pressure changes associated with these three choices of the strain rate. Panel (c), Blue data show the first peak in the derivative of the precipitated fraction, $df/dt$ versus strain rate. This corresponds to the nucleation of precipitates on the grain boundary. Beyond a critical strain rate of $\dot{\epsilon}=3\times 10^{-6}$ dt-1, the second peak in $df/dt$ appears, shown here in red.
@@
Upon further increase in $\dot{\epsilon}$, the deformation of the system causes the GB to buckle as shown in Figure 2. This buckling further provides preferential sites of precipitate nucleation due to a shifted contact angle as well as curvature induced solute segregation at the cusps. Furthermore, the formation of these precipitates relieves the strain energy stored within these regions. As shown in Figure 3, the resulting pressure for high value of $\dot{\epsilon}$ exhibits only a single prominent peak, which accounts for both a buildup through this buckling mechanism and then a rapid release due to precipitation.

### v2

# Deformation assisted precipitation in binary alloys: A competition of time-scales

@@ Abstract
We consider the process of precipitation in binary alloys in the presence of mechanical deformation. It is commonly observed that mechanical deformation prior to or during precipitation leads to microstructure with excess defects, which allows for enhanced precipitate nucleation and growth rates [1, 2, 3]. To investigate this phenomenon, we employ a two-dimensional phase-field crystal alloy model endowed with a temperature dependent mobility, making it capable of recovering isothermal transformation (TTT) diagrams with a characteristic inflection point (nose) about a critical temperature. We examine the variation in the time-scale of precipitation and its connection to the time-scale of the applied deformation, focusing on the roles of atomic defects in the processes involved. Our results indicate that precipitation is initially delayed through application of a deformation until a critical strain is achieved, beyond which precipitation proceeds more rapidly, assisted by plastic deformation such as grain boundary serration or dislocation nucleation. We show that the evolution of the precipitated fraction, $f(t)$, departs from classical Avrami behaviour. Specifically, $df/dt$ develops two peaks indicative of a “plateau"-like inflection in $f(t)$, signalling the transition to defect assisted precipitate nucleation. We analyze these plateaus as a function of the deformation rate and demonstrate that the they exhibit a discontinuous bifurcation as the time-scale of applied deformation is increased. These findings are compared to and found to be consistent with experiments.
@@
I Introduction
Mechanical deformations driving out-of-equilibrium behaviours are ubiquitous in current and emerging material processing techniques such as additive manufacturing, laser shock peening, friction stir-welding, etc. [4, 5, 6].
The material conditions left behind in the wake of these processes provide inflated driving forces for diffusive and phononic relaxation due to increased saturation of solid solutions and the presence of point defects. This is in sharp contrast to classical casting conditions [7, 8]. An emphasis on the physics governing the aforementioned material processes has thus been of interest to the materials science community, particularly how such processes deviate from classically understood mechanisms of precipitation, recrystallization and recovery in metal alloys.
Precipitation readily occurs at sites of higher energy, such as defects and Grain Boundaries (GBs) [9]. Mechanical deformation of crystalline solids produces many such sites, thus modifying the nucleation, growth, and coarsening properties encountered [3].
Precipitation can thus be seen as competitive with other forms of strain relaxation eg. recrystallization and recovery in metals and their alloys. These processes all use stored driving forces to relieve stress accruing at high energy sites within a deformed sample. Indeed, thus has been studied previously [1, 2], where experiments and numerical simulations have examined modifications to isothermal transformation (TTT) diagrams in the presence of mechanical deformation. Analytical models have likewise been formulated to quantify this interactive growth process [10]. These models typically require a priori knowledge of specific material properties which are difficult to extract such as activation energies and defect fractions. Furthermore, these models require knowledge of the nucleation and growth mechanisms of both recrystallized and precipitated grains.
@@ I Introduction
This paper is structured as follows: In Section II, we introduce the details of the XPFC model we employ including temperature dependent mobilities. In Section III, we describe the numerical approach we take to model a precipitating system under continuous deformation. In section IV, we present our results, including the demonstration of a second, deformation induced time-scale in the dynamics of precipitation. Finally in Section V, we briefly provide a future outlook and describe how our results may be interpreted.
II Phase-Field Crystal Model
@@
II.1 Free Energy Functional
In order to investigate the atomistic behaviour of precipitation in the presence of deformation, this work examines the prototypical scenario of an equal tilt bicrystal undergoing a continuous biaxial deformation with strain rate parameter $\dot{\epsilon}$, which shall be clarified later. To this end, we employ a structural phase field crystal (XPFC) model with a free energy functional that models substitutional binary alloys and has previously been applied to studies involving precipitation [23, 24], as well as deformation in such systems [25]. Our free energy is formulated as follows,
@@ II.1 Free Energy Functional
where $c_{0}$ is the concentration at the reference. The final, excess, term of the free energy controls the length scale and strength of atomic-scale interactions. The interactions are embedded in the two-point density-density correlation function $C_{nn}$ which interpolates between pure $A$ and pure $B$ structures through compositionally dependent functions $\chi_{A}(c)=1-3c^{2}+2c^{3}$, $\chi_{B}(c)=1-3(1-c)^{2}+2(1-c)^{3}$ such that $C_{nn}^{(2)}(\mathbf{r}-\mathbf{r^{\prime}})=\chi_{A}C_{AA}^{(2)}(\mathbf{r}-\mathbf{r^{\prime}})+\chi_{B}C_{BB}^{(2)}(\mathbf{r}-\mathbf{r^{\prime}})$. In XPFC, the pure species correlation functions, $C_{ii}^{(2)}$ are defined in Fourier space (with wave-vector magnitude $k$) by Gaussian functions centered about the dominant reciprocal lattice wave-vector ($k=k_{i}$) and $k=0$ modes of each species respectively,
@@
where $i=A,B$. The parameter $T_{i}^{0}$ controls the temperature scaling of the corresponding correlation function, $\sigma_{i,0}$ and $\sigma_{i,1}$ control the Gaussian widths of the $k=0$ and dominant wave-vector respectively. Finally, $A_{i,0}$ is a constant that controls the amplitude of the $k=0$ mode. In this letter, we work in two spatial dimensions and employ a single wave-vector in our correlation to stabilize a hexagonal structure in each solid phase. The details of the parameters we employ and the resulting phase-diagram are shown in Appendix A.
II.2 Dynamics
@@
where $\mu_{n/c}$ are the chemical potentials, $M_{n/c}$ are the mobilities and $\xi_{n/c}$ are noise sources applied to the $n$ and $c$ fields respectively. The noise terms satisfy the fluctuation dissipation relations, $\langle\xi_{n/c}(\textbf{r},t)\xi_{n/c}(\textbf{r'},t^{\prime})\rangle=N_{n/c}M_{n/c}k_{B}T\delta(\textbf{r}-\textbf{r'})\delta(t-t^{\prime})$ where $N_{n/c}$ are model parameters subsuming remaining constant prefactors. The noise terms here are implemented to allow for stochastic variations between different realizations of the same initial conditions. Without such fluctuations, the driving force for precipitation presented by enhanced solute segregation to dislocations and GBs may not be sufficient to nucleate precipitates despite a quench into the metastable region. In such systems, fluctuations act to drive the concentration field over the barrier present.
We also add an additional time-scale through the parameter $\beta$ for quasi-phonon relaxation into the density equation. Here, $\beta$ defines the effective length-scale over which quasiphonons propagate undamped[20]. A two-time scale coupling to $c$ is not examined here as we do not anticipate any transport of solute on time-scales commensurate with elastic relaxation.
It is worthwhile to note that Eqs. (4) and (5) are not strictly consistent in the regime where the system’s average density has significant variations [26]. A more rigorous treatment of the dynamics has been implemented in [27]. However, as the common tangents between the free energy wells of our system exist at small and nearly identical average densities, this approximation is justified.
We expand on the model by introducing temperature dependent mobilities into Eqs. (4) and (5). This allows for the generation of TTT diagrams that can be controlled through model parameters. An Arrhenius form is assumed for the mobilities of both $n$ and $c$,
@@
where $E_{A}^{n/c}$ and $M_{n/c}^{0}$ are the activation energies, and the mobilities at $T=0$ respectively, which control the curvature and position of the inflection point of a given isothermal transformation (TTT) curve for precipitation. Formally, the mobilities $M_{n/c}$ are non-linearly coupled to the mobilities of the individual atomic species. In this paper, we treat them as mutually independent model parameters. It is also noted that the modified mobilities will shift the quasiphonon propagation decay length. No qualitative differences are observed in our results due to such shifts.
@@
Figure 1: The TTT diagram of precipitation at zero strain rate. The three sets of curves represent different fractions of completed transformation. A characteristic inflection or nose shape can be observed over the range of temperatures selected as a result of the temperature dependent mobilities defined in Eq. (6). Curves are plotted for a selection of $E_{A}^{c}$, which monotonically controls the reaction rate as shown.
@@
Figure 2: The sequence of events of a simulation of strain enhanced precipitation. Panels ($\text{a}_{1}$)-($\text{a}_{4}$) show the evolution of the composition field and panels ($\text{b}_{1}$)-($\text{b}_{4}$) show the behaviour of the local orientation field. Panels ($\text{c}_{1}$)-($\text{c}_{4}$) show a zoom-in of the density field in the highlighted magenta regions of panel rows (a) and (b). The times and strains along each axis associated with each column of panel rows (a), (b) and (c) are shown at the bottom of the figure. Panel (d) shows the associated evolution of the precipitated fraction $f$ as well as $df/dt$ in black and red, respectively, with the values corresponding to panels ($\text{a}_{1}$)-($\text{a}_{4}$) labelled with circles. Panel (e) shows the histogram of orientations of panels ($\text{b}_{1}$) and ($\text{b}_{4}$). Panel (f) shows a schematic of our system including the directions of tension and compression. Our tilt axis is parallel to the y axis and we consider an axisymetric tilt where $\theta_{1}\approx-6^{\circ}$ and $\theta_{2}\approx+6^{\circ}$. The first peak in ${df}/{dt}$, corresponding to panels ($\text{a}_{2}$), ($\text{b}_{2}$) is associated with the precipitation on the serrated interface. The second peak, corresponding to panels ($\text{a}_{3}$), ($\text{b}_{3}$) coincides with precipitate nucleation on dislocations. Panels ($\text{b}_{4}$) and ($\text{c}_{4}$) show the late stage coarsening behaviour of the system, which includes the reorientation of A-rich regions close to the grain boundary down to orientations intermediate between zero and the initial orientation of ±$\pm 6^{\circ}$.
III Numerical Simulation Details
@@
For all simulations described, we performed a quench below the solvus line into solid-solid coexistence. Full simulation parameters are presented in the Supplementary Material. We initialize the density field of the crystalline regions by seeding in an equal tilt bicrystal through a one-mode expansion. The lattice constant of this solid solution is set to be commensurate to the value attained by the interpolated correlation function for the given choice of supersaturation. The system is set to be initially strain-free by performing a minimization over the angle and number of atoms in the x and y directions for one (and by symmetry the other) grain as in [28]. The initial misorientation during this minimization was kept within the range 12.00±$\pm 0.05^{\circ}$. The average density and supersaturation are set to be initially uniform and equal to 0.05 and 0.15, respectively.
The density field is next allowed to relax by setting $M_{c}$ to zero. This relaxation entails the evolution of only Eq. (4) until the amplitude of the density field reaches an equilibrated value in the bulks of the grains. After this time, Eq. (5) is solved concomitantly with Eq. (4) introducing solute dynamics. At this stage, an isochoric biaxial deformation is applied to the system. This is done by varying the system discretization along both axes, similarly to the approach employed in [25]. We note that the strain rate parameter $\dot{\epsilon}$ is higher than the true strain rate by approximately two orders of magnitude due to the numerical methodology employed for deformation. In the proceeding discussion, for convenience, we will nevertheless refer to $\dot{\epsilon}$ as the strain rate. We provide further numerical details in Appendix B.
@@
Over the course of simulations we track the grand potential energy (analogously, the dimensionless system pressure $\bar{P}=P/k_{B}T^{0}\rho^{0}V$ [29]), precipitated fraction $f$, and local orientation. The orientation is calculated as in [30] and $f$, is calculated through thresholding in composition in conjunction with a Fourier space mask sharply peaked around the dominant wave-vector of the precipitating phase.
IV Results & Discussion
@@
We first examine the situation where $\dot{\epsilon}=0$. While precipitation has been studied with XPFC in the past, to our knowledge no study has been performed with temperature dependent mobilities as in Eq. (6). Over the course of each isothermal simulation, fluctuations drive the formation of precipitates along the GB of the system. Over all temperatures and activation energies examined, the precipitate fraction obeys an Avrami-like behaviour
@@
where $m$ is an exponent determined by the dimensionality and mechanism of the transformation, $f^{eq}$ is the equilibrium precipitate fraction and $\tau_{0.5}$ is a characteristic time-scale. The system pressure was found to exhibit a single dominant peak corresponding to the end of precipitate nucleation. Beyond this time, the system reduces its free energy and pressure through diffusion of solute out of the surrounding supersaturated bulk and into precipitate growth.
Figure 1 shows the TTT curves for a range of model temperatures, demonstrating the inflection about a critical temperature. The scaling of this surface is controlled by the usual driving force for precipitation (competition between bulk and surface energies) as well as the new mobility used here, which also affects the noise strength. At each temperature, the bulk energies are determined through the curvatures of the free energy landscape about the equilibrium wells of Eq. (1) in the space of average density, supersaturation and amplitude of the atomic density field $n$. The surface energy will be dependent on all of the above along with the effective gradient coefficient of the amplitude arising from the excess energy term, as well as $W_{c}$. These parameters are all held fixed for each value of $T$ along the curves shown in Figure 1.
@@
The evolution of a precipitating system under mechanical deformation is shown in Figure 2. At lower values of $\dot{\epsilon}$, $f$ obeys the same qualitative Avrami behavior as the undeformed case. Prior to precipitation, in order to compensate for the applied deformation, the density field undergoes elastic accommodation. The characteristic time-scale, $\tau_{0.5}$, for the precipitation is observed to initially increase with $\dot{\epsilon}$. This occurs due to an accelerated rotation of the bicrystal with increasing $\dot{\epsilon}$, causing a reduction in the effective driving force for precipitation along the boundary. The increase we observe in $\tau_{0.5}$ follows a simple $\exp(-\Delta\gamma^{2})$ dependence in the time-scale for precipitation. Here $\Delta\gamma=\gamma_{f}-\gamma_{i}$ where $\gamma_{f}$ is the precipitate-matrix surface energy and $\gamma_{i}$ is the initial surface energy, whose misorientation dependence has been shown to reproduce the correct Read-Shockley scaling [31]. We have also performed simulations in the undeformed case with varying orientation, recovering the same Read-Shockley scaling.
@@
Figure 3 plots the system pressure as a function of time at different strain rates $\dot{\epsilon}$. As in the zero deformation case, at low $\dot{\epsilon}$ the pressure is first found to exhibit a single dominant peak at the initial stages of precipitation, followed by oscillations indicating accommodations by the GB and the formation of any dislocations necessary to support the growth of the population of precipitates.
@@
Figure 3: The system pressure is plotted as a function of system time for different choices of the strain rate. At lower values of $\dot{\epsilon}$ (dark red), an initial peak followed by oscillations indicating nucleation of dislocations post-precipitation is observed. At larger $\dot{\epsilon}$ (orange-yellow) a second peak emerges corresponding to a new dislocation assisted precipitation mechanism. At the largest range of $\dot{\epsilon}$ simulated (blue-pink), a single peak emerges again indicating precipitation assisted through GB serration.
@@
As $\dot{\epsilon}$ is increased further beyond $2\times 10^{-6}$, we observe three mechanisms that contribute to a systematic deviation in the precipitation mechanism, each activating in order of increasing $\dot{\epsilon}$.
Figure 4: The times when $f/f^{eq}$ attains $0.05,0.50$ and $0.95$ plotted against strain-rate. Also shown in blue is the time at which the system achieves its maximum in pressure $t_{P(max)}$ at each strain-rate. At low $\dot{\epsilon}$, $t_{P(max)}$ corresponds to the initialization of precipitation. At a critical $\dot{\epsilon}=2\times 10^{-6}$, $t_{P(max)}$ diverges significantly from its value during the initiation of precipitation. This occurs because the prominence of the second, plasticity induced, pressure peak surpasses the peak associated with precipitation, as shown in Figure 3. Upon further increase of $\dot{\epsilon}$ beyond $5\times 10^{-6}$, serration of the GB occurs and $t_{P(max)}$ monotonically decreases past its initial value.
@@
At the lower range of $\dot{\epsilon}$, we observe nucleation of dislocations from initial precipitates or the GB, also shown in Figure 2. It is well known that dislocations act as preferential nucleation sites for precipitation [9]. This has also been examined numerically in prior studies using the XPFC model [23, 24]. Excess solute may diffuse to not only the GB but also towards the dislocations flowing into the bulk of the system, inducing further precipitation at sites away from the GB. Dislocation nucleation events are characterized by a new peak in the system pressure that progressively increases in prominence and occurs after initial precipitation along the GB. This peak is subsumed into the pressure shift due to precipitation at intermediate $\dot{\epsilon}$ as shown in Figures 3 and 4. The same effect also leads to the emergence of a second peak in $df/dt$, shown in Figure 5, corresponding to additional precipitate nucleation mechanism.
@@
Figure 5: Panel (a) shows $f$ (green) and $df/dt$ (black), indicating the first and second peaks of $df/dt$ (blue and red respectively). The data are plotted for three values of progressively larger $\dot{\epsilon}$. The low $\dot{\epsilon}$ scenario demonstrates a single peak whereas the second and third cases exhibit a prominent second peak in $df/dt$, associated with dislocation nucleation. Panel (b) shows the pressure changes associated with these three choices of $\dot{\epsilon}$. In panel (c), blue data show the first peak in $df/dt$ vs. $\dot{\epsilon}$. This corresponds to the nucleation of precipitates on the GB. Beyond a critical strain rate of $\dot{\epsilon}=3\times 10^{-6}$ dt-1, the second peak in $df/dt$ appears, shown here in red.
@@
Upon further increase in $\dot{\epsilon}$, deformation causes the GB to serrate as shown in Figure 2. This serration provides preferential sites of precipitate nucleation due to shifted contact angle and curvature induced solute segregation. Furthermore, formation of these precipitates relieves the strain energy stored in the serrated GB. As shown in Figure 3, the resulting pressure for high values of $\dot{\epsilon}$ exhibits a single prominent peak, accounting for both a buildup through this serration mechanism and subsequent release from precipitation.